OpenStudy (anonymous):

Write the slope intercept equation for the line with slope -3/7 and y- intercept (0,-5) and graph the line. Show all work please.

OpenStudy (anonymous):

y-y1 = m(x-x1) y--5 = -3/7 (x-0) y+5 = -3x/7 y=-3x/7 -5